Krypton-87 has rate constant of 1.5 x 10 -4 / s. What is the activity of a 2.0 mg sample?


1
Expert's answer
2021-05-26T07:08:18-0400

Solution:

Krypton-87 (86.91 g/mol)

Convert the 2.00 mg sample of krypton-87 into numbers of atoms:

(2.00 mg Kr-87)(1g / 1000 mg)(1 mol Kr-87 / 86.91 g)(6.022×1023 atoms/mol) = 1.3858×1019 atoms


Use the rate equation to calculate the activity:

r = (1.5×10-4 s-1)(1.3858×1019 atoms) = 2.0787×1015 atoms decay s-1


Convert atoms s-1 into curies (Ci):

(2.0787×1015 atoms s-1)(1 Ci / 3.700×1010 atoms s-1) = 56181.1 Ci


Answer: 56181.1 curies